shift summary, day 08/04/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Wed, 4 Aug 1999 15:30:36 -0500

Shift Summary for day shift, 08/04/99

On Shift: Ashkan Alavi, Andera Soddu

Hours of beam available at all intensities: 8.0/8.5
Hours of beam available at nominal intensity: 7.0/8.5
Hours of E832 data (special runs not included): 8.0/8.5
Hours of special runs (laser scans, muon runs...):

Accounting of lost time: Main Injector power supply problem, adjusting
high intensity beam.

========================================================================
Runs ended this shift:
Run 14103 - 9.8M events - BEAM_832
Run 14104 - 8.9M events - BEAM_832

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: smooth E832 datataking after setting the absorber back
in the right position early in the shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14103 07:00 Inherit run from the exhausted owl shiftees.

07:14 Serious Alarm keeps complaining about high rates
for BEAM01 and BEAM04. Also DC1 & 2 are drawing
too much current. Arun correctly suspects the
absorber is still in the wrong position, even
though the read out is correct. Rick Ford moves it
from stop 1 to 2 and back to 1. Now everything
looks fine!

07:20 Stop the run.

14104 07:30 Start the new run.

07:35 Onsplit requests tape QKA006 be mounted on tps2d4
ktev3.

07:40 Noticed the CsI online was dead. fix_disp fixes.

08:40 level 3 online kumac shows one the monitoring
process on ktev3 is not running. Ed doesn't think
it would affect the data for now.

10:15 On Rick Ford's request, AA disabled sdaq
monitoring of NM1U magnet which turns the epicure
button red often, until the read out is
fixed. So please don't ignore epicure errors on
the main slow da screen.

10:28 Onsplit requests tape QKSM09 be mounted on tps3d2
ktev4.

11:20 Call MCR to request high intensity 6.8E12.

11:25 MCR calls back. They are starting to raise the
intensity. It could happen in 30sec or 3hrs!

179 11:27 Stop the run. Do a drt_fresh_start.

11:38 Onsplit requests tape QKLA11 be mounted on tps2d3
ktev3.

11:40 MI power supply problem. Experts investigating.

12:05 Beam is back but with low intensity.

14105 12:10 Start the run, hoping to get high intensity.

12:17 MCR calls. They can not inject beam to the TeV
with our requested intensity. Experts are working
on it. Continue with low intensity for now.

13:00 DN connects CO2 dewar to gas system instead of
bottles. Pressure is to be read at the regulator
attached to the dewar now. A back up bottle will
be installed later. He also ordered 8 new bottles
of Ar-Ethane for delivery tomorrow.

39 13:05 First high intensity spill.

14:08 Onsplit requests tape QKNL17 be mounted on tps3d2
ktev2.

14:15 Occasionally the SEM goes up to 7.2E12, asked MCR
to keep it below 7E12, otherwise stable.

========================================================================
Follow up Notes on Previous Shift:
Continue taking E832 data.

========================================================================
Items to Follow up on Next Shift: